# Singularity resolution and inflation from an infinite tower of regularized curvature corrections

### v1

@@ Abstract
$\begin{array}[]{|c|c|c|c|c|}\hline\cr c_{n}&\text{Horndeski Functions}&F(x)&H^{2}&f(r)\\ \hline\cr\hline\cr\frac{1-(-1)^{n}}{2n}&\begin{aligned} &G_{2}=-2\Lambda+\frac{4X\left(28\ell^{4}X^{2}-3\right)}{(1-4\ell^{4}X^{2})^{2}}+\frac{6\tanh^{-1}\left(2\ell^{2}X\right)}{\ell^{2}}\\ &G_{3}=\frac{2\left(1-20\ell^{4}X^{2}\right)}{\left(1-4\ell^{4}X^{2}\right)^{2}}\\ &G_{4}=\left(1-4\ell^{4}X^{2}\right)^{-1}\\ &G_{5}=-4\ell^{2}\left[\frac{\ell^{2}X}{1-4\ell^{4}X^{2}}+\frac{1}{2}\tanh^{-1}\left(2\ell^{2}X\right)\right]\end{aligned}&\frac{\tanh^{-1}\left(\ell^{2}x\right)}{\ell^{2}}&\frac{\tanh\left[8\pi\ell^{2}\rho/3\right]}{\ell^{2}}&\frac{r^{2}}{\ell^{2}}\tanh\left[\frac{\ell^{2}}{r^{2}}f_{\rm GR}\right]\\ \hline\cr 1&\begin{aligned} &G_{2}=-2\Lambda+\frac{8\ell^{2}X^{2}\left(1+6\ell^{2}X\right)}{\left(1-2\ell^{2}X\right)^{3}}\\ &G_{3}=\frac{2\left(1-10\ell^{2}X\right)}{\left(1-2\ell^{2}X\right)^{3}}\\ &G_{4}=\left(1-2\ell^{2}X\right)^{-2}\\ &G_{5}=-4\ell^{2}\left(\frac{1-2\ell^{4}X^{2}}{\left(1-2\ell^{2}X\right)^{2}}+\log\left(\frac{2\ell^{2}X}{1-2\ell^{2}X}\right)\right)\end{aligned}&\frac{x}{1-\ell^{2}x}&\frac{8\pi\rho}{3+\ell^{2}8\pi\rho}&\frac{r^{2}f_{\rm GR}}{r^{2}-\ell^{2}f_{\rm GR}}\\ \hline\cr\begin{aligned} &\frac{\Gamma(n/m)\delta_{0,k}}{\Gamma(1/m)\Gamma((n+m-1)/m)},\\ &k=(n-1)\bmod m\end{aligned}&\begin{aligned} &G_{2}=-2\Lambda+\frac{2^{m+2}\ell^{2m}X^{m+1}\left(2m-1+3\ell^{2m}X^{m}2^{m}\right)}{(1-2^{m}\ell^{2m}X^{m})^{(2m+1)/m}}\\ &G_{3}=2\frac{1-2^{m}(3+2m)\ell^{2m}X^{m}}{(1-2^{m}\ell^{2m}X^{m})^{(2m+1)/m}}\\ &G_{4}=\left(1-2^{m}\ell^{2m}X^{m}\right)^{-(m+1)/m}\\ &G_{5}=-\int\frac{2^{m}(m+1)\ell^{2m}X^{m-2}}{(1-2^{m}\ell^{2m}X^{m})^{(2m+1)/m}}\mathrm{d}X\end{aligned}&\frac{x}{\left(1-\ell^{2m}x^{m}\right)^{1/m}}&\frac{8\pi\rho/3}{\left(1+\left(\frac{8\pi\rho}{3}\ell^{2}\right)^{m}\right)^{\frac{1}{m}}}&\frac{f_{\rm GR}}{\left(1+\left[-\frac{\ell^{2}}{r^{2}}f_{\rm GR}\right]^{m}\right)^{\frac{1}{m}}}\\ \hline\cr 1/n&\begin{aligned} &G_{2}=-2\Lambda+\frac{4X\left(10\ell^{2}X-3\right)}{\left(1-2\ell^{2}X\right)^{2}}-\frac{6\log\left(1-2\ell^{2}X\right)}{\ell^{2}}\\ &G_{3}=\frac{2\left(1-6\ell^{2}X\right)}{\left(1-2\ell^{2}X\right)^{2}}\\ &G_{4}=(1-2\ell^{2}X)^{-1}\\ &G_{5}=-2\ell^{2}\left(\frac{1}{1-2\ell^{2}X}-2\tanh^{-1}\left(1-4\ell^{2}X\right)\right)\end{aligned}&-\frac{\log\left(1-\ell^{2}x\right)}{\ell^{2}}&\frac{\left(1-e^{-8\pi\rho\ell^{2}/3}\right)}{\ell^{2}}&\frac{r^{2}}{\ell^{2}}\left(e^{\frac{\ell^{2}}{r^{2}}f_{\rm GR}}-1\right)\\ \hline\cr\end{array}$
@@
Supplemental Material B Scalar field current for black holes with $k\neq 0$

### v2

@@ Abstract
[…]
@@
Supplemental Material B Scalar field current for black holes with $k\neq 0$, and Carter-Penrose diagram for a regular black hole
@@
In Fig. 2 we present the Carter-Penrose diagram of the black hole solution (15). This diagram is generic for the geodesically complete regular black holes obtained in this framework.
[…]
Figure 2: Carter-Penrose diagram of the black hole solution in Eq. (15). Region $\rm I$ corresponds to an asymptotically AdS region; Region $\rm II$ corresponds to the interior of the black hole; Region $\rm III$ corresponds to an asymptotically AdS region in a parallel universe.
Supplemental Material C Generalized Friedmann equations for a spatially-curved FLRW universe
In the main text we focused on the spatially-flat FLRW case. Here, we generalize our results for the spatially-curved FLRW case, with curvature $k\in\{-1,0,1\}$, whose metric is
$\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a(t)^{2}\left(\frac{\mathrm{d}r^{2}}{1-kr^{2}}+r^{2}\mathrm{d}\theta^{2}+r^{2}\sin^{2}\theta\mathrm{d}\varphi^{2}\right).$ (SM.3)
Using the same scalar field profile as in the spatially-flat case, Eq. (7), we find that the equation of motion for the scalar is identically solved, and we obtain the generalized Friedmann equations
$\frac{1}{\ell^{2}}\sum_{n=1}^{\infty}c_{n}\left(\ell^{2}H^{2}\right)\left[1+n\frac{k}{a^{2}H^{2}}\right]=\frac{8\pi}{3}\rho.$ (SM.4)
As an example, for the case studied in the main text, $c_{n}=(1-(-1)^{n})/(2n)$, we obtain
$\frac{\tanh^{-1}\left(\ell^{2}H^{2}\right)}{\ell^{2}}+\frac{k}{\left(1-\ell^{4}H^{4}\right)a^{2}}=\frac{8\pi}{3}\rho,$ (SM.5)
and for $c_{n}=1$ we get
$\frac{H^{2}}{1-\ell^{2}H^{2}}+\frac{k}{(1-\ell^{2}H^{2})^{2}a^{2}}=\frac{8\pi}{3}\rho.$ (SM.6)
Taking the continuity equation into account, from which we find $\rho\sim a^{-3(1+\omega)}$, or equivalently , $a\sim\rho^{-\frac{1}{3(1+\omega)}}$, it is possible to show that even in the spatially-curved FLRW case, these equations limit $H^{2}<1/\ell^{2}$, and therefore the curvature is bounded.